OpenStudy (anonymous):

Medal Reward!! Customers of a phone company can choose between two service plans for long distance calls. The first plan has a $19.00 monthly fee and charges an additional $0.13 for each minute of calls. The second plan has a $9.00 monthly fee and charges an additional $0.18 for each minute of calls. For how many minutes of calls will the costs of the two plans be equal?

OpenStudy (anonymous):

19+0.13m=9+0.18m 10=0.5m 20=mins